OpenStudy (anonymous):

square binomial of x2 + 12x + 36

OpenStudy (anonymous):

What is 6*6 = ?? And 6 + 6 = ??

OpenStudy (anonymous):

\[x^2 + 12x + 36 \implies x^2 + 6x + 6x + 36\] Can you know factor it out or not??

OpenStudy (anonymous):

i like to know how to factor it

OpenStudy (dumbcow):

its a perfect square right (x+a)^2 = (x+a)(x+a) = x^2 +ax +ax+a^2 --> a^2 = 36 --> a = 6

OpenStudy (anonymous):

I give you the example: Suppose: We are given with: \(x^2 + 2x + 2x + 4\) So, group the terms first: \((x^2 + 2x) + (2x + 4)\) = \(x(x+2) + 2(x+2)\) So, \((x+2)(x+2)\)

OpenStudy (anonymous):

In place of 2 there is 6 and in place of 4 there is 36..

OpenStudy (theviper):

Oh I m sorry I gave full answer :(

OpenStudy (anonymous):

that's ok

OpenStudy (anonymous):

It is better to delete the post if you have given the answer than to say sorry..

OpenStudy (anonymous):

Thanks.. @TheViper

OpenStudy (theviper):

yes I have already deleted but my internet is slow :D